Fakeland has 2.3 million employed people and 0.2 million unemployed people. The size of the labour force is fixed. The job finding rate is 22% and the separation rate is 3%.

(a) What is Fakeland’s current unemployment rate?

(b) How many unemployed workers will find a job this period?

(c) How many employed workers will lose their job this period?

(d) Given your answers above, will the number of unemployed people increase or decrease over time?

(e) Compute Fakeland’s steady state unemployment rate. (f) In the steady state, how many unemployed workers find a job in each period? How many employed workers lose their job in each period?
Answer #1

A) labor force = employed + unemployed = (2.3 + 0.2) millions = 2.5 millions

Unemployment rate = (no. Of unemployed/labor force)*100 = (0.2 million/2.5 millions)*100 = 8%

B) 22% of unemployed workers will find job this period, that is, (0.2 * 22%) million = 0.044 million people will be employed.

C) (2.3 * 3%) million = 0.069 million people will loose their job.

D) Now the number of unemployed people = (0.2 - 0.044 + 0.069) million = 0.225 million. It means no. Of unemployed people increased over time.

E) Steady state unemployment rate = s/(s + f)

Here, job separation rate (s) = 3% = 0.03 and job finding rate (f) = 22% = 0.22

Steady state unemployment rate = 0.03/(0.03 + 0.22) = 0.03/0.25 = 0.12 = 12%